OpenStudy (anonymous):

The equation of a line is y=-2/3x-1. What is the equation of the line that is perpendicular to the first line and passes through (–4, 3)?

OpenStudy (anonymous):

a) y=3/2x+9 b) y=2/3x+17/3 c) y=3/2x-3 d) y=-3/2x-3